OpenStudy (anonymous):

An electronics manufacturer recently created a new version of a popular device. It also created this function to represent the profit, P(x), in tens of thousands of dollars, that the company will earn based on manufacturing x thousand devices: P(x) = -0.16x2 + 21.6x – 400.

OpenStudy (anonymous):

a. The profit function for the first version of the device was very similar to the profit function for the new version. As a matter of fact, the profit function for the first version is a transformation of the profit function for the new version. For the value x = 40, the original profit function is half the size of the new profit function. Write two function transformations in terms of P(x) that could represent the original profit function.

OpenStudy (anonymous):

medal me and ill help

rebeccaxhawaii (rebeccaxhawaii):

you don't get a medal for not helping -.- @jchris25

OpenStudy (dumbcow):

Let f(x) be original profit function 1 transformation could be a vertical shift \[f(x) = P(x) + k\] Given: \[f(40) = \frac{1}{2}P(40)\] \[\rightarrow P(40) +k = \frac{1}{2} P(40)\] \[k = -\frac{1}{2} P(40)\]